Dhaka University took a test and the final grades have a mean of 100 and a standard deviation of 11. If we can approximate the distribution of these grades by a normal distribution, what percent of the students

a) scored higher than 120?

b) should pass the test (grades≥90)?

c) should fail the test (grades<90)?


1
Expert's answer
2021-05-02T16:01:50-0400

Denote

"\overline{x}=100"

"s=11"


a) "P(X>120)=P(120<X<+\infty)="

"F(+\infty)-F\left(\dfrac{120-\overline{x}}{s}\right)="

"1-F\left(\dfrac{120-100}{11}\right)=" "1-F\left(\dfrac{20}{11}\right)\approx" "1-F\left(1.8182\right)\approx1-0.9655\approx0.0345"

Percent of the students scored higher than 120 is "0.0345\times100=3.45\%"


b) "P(X\geq90)=P(90\leq X<+\infty)="

"F(+\infty)-F\left(\dfrac{90-\overline{x}}{s}\right)="

"1-F\left(\dfrac{90-100}{11}\right)=" "1-F\left(-\dfrac{10}{11}\right)\approx"

"1-F\left(-0.9091\right)\approx1-0.1817\approx0.8183"

Percent of the students should pass the test is "0.8183\times100=81.83\%"


c) "P(X<90)=P(-\infty<X<90)="

"F\left(\dfrac{90-\overline{x}}{s}\right)-F(-\infty)="

"F\left(\dfrac{90-100}{11}\right)-0=F\left(-\dfrac{10}{11}\right)\approx" "F(-0.9091)\approx0.1817"

Percent of the students should fail the test is "0.1817\times100=18.17\%"


Note that b) and c) must satisfy "81.83\%+18.17\%=100\%"